# Agile Tradespace Exploration for Space Rendezvous Mission Design via Transformers

### v1

@@ 5 Spacecraft Rendezvous in perturbed Keplerian orbits
This paper focuses on mission design for on-orbit servicing operations targeting multiple resident space objects in Earth orbit. In this scenario, a servicing spacecraft (servicer) executes an approach maneuver to a predefined waypoint in the vicinity of a target spacecraft (target). During the preliminary trajectory design phase of such missions, it is essential to conduct rapid trade studies between time of flight and total control cost across a diverse set of candidate targets, each characterized by distinct orbital elements and, consequently, unique relative dynamics.
@@ 5.1 Optimal Control Problem Formulation
The trajectory optimization problem described above can be formulated as a discrete-time MO-OCP for chance-constrained, fault-tolerant rendezvous trajectory design [18] as follows:
@@
The state $\boldsymbol{x}\in\mathbb{R}^{6}$ is expressed in Relative Orbital Elements (ROE) [10][57]. Let the Keplerian orbital elements be $\boldsymbol{\alpha}_{c}=[a,e,i,\Omega,\omega,M]$, where $a$ is the semi-major axis, $e$ the eccentricity, $i$ the inclination, $\Omega$ the right ascension of the ascending node, $\omega$ the argument of periapsis, and $M$ the mean anomaly. The ROE are a particular set of differences between the orbital elements of the target and the servicer (also called the deputy), denoted as $\boldsymbol{\alpha}$ and $\boldsymbol{\alpha}_{d}$, respectively. The objective is to simultaneously minimize the total control cost—modeled as a sum of impulsive maneuver magnitudes, and the terminal time represented by the discrete horizon length $N$, leading to a time of flight of $N\Delta t$. Eq. (5c) describes the discrete-time ROE dynamics with impulsive velocity change $\boldsymbol{u}\in\mathbb{R}^{3}$ applied in the Radial/Tangential/Normal (RTN) frame. The State Transition Matrix (STM) $\Phi(t_{k+1},t_{k})\in\mathbb{R}^{6\times 6}$ and control input matrix $\Gamma_{k}=\Gamma(t_{k})\in\mathbb{R}^{6\times 3}$ are expressed as an analytical function of the target’s orbital elements but not the relative states themselves [10][57], which offers a convenient expression as a linear time-varying system. Additionally, Eqs. (5e) and (5f) represent boundary conditions and maximum instantaneous control magnitude constraints, respectively. Finally, Eq. (5d) indicates the passive safety constraint, which ensures the deputy is outside of the ellipsoidal safety domains around the target not only for the entire duration of the controlled trajectory $t\in[0,t_{f}]$ but also for a predefined duration ($\tau\in[0,\tau^{s}]$) of the drift trajectories after the sudden loss of control authority. The geometry of the safety domains at epoch $t_{k}$ is defined using the diagonal matrix $P_{k}\in\mathbb{R}^{6\times 6}$, which encodes the semi-axes of the ellipsoid within the RTN frame. In this case study, two safety domains are considered: the Approach Ellipsoid (AE) and the Keep-Out Zone (KOZ), both modeled as spheres. At the switching epoch, $t_{\text{switch}}$, the active safety domain transitions from the larger AE to the smaller KOZ, reflecting the practical scenario in which the allowable safety region contracts as the chaser spacecraft closes in on the target. Matrix $\Psi_{k}=\Psi(t_{k})\in\mathbb{R}^{6\times 6}$ denotes the first-order linear mapping between the ROE and the matrix expression of the relative position and velocity vector in the Cartesian coordinates within the RTN frame (denoted as $\boldsymbol{x}_{\text{RTN}}$) [10]. The passive safety constraint is formulated as a chance constraint to account for the inherent uncertainty that comes from navigation, control execution error, and unmodeled acceleration. At each epoch, the risk factor (i.e., failure probability) $\Delta_{k}$ is chosen as a parameter, and the probabilistic satisfaction of the constraint is enforced. Assuming Gaussian state uncertainty, the drifted mean state and covariance propagated from the time immediately after the control epoch $t_{k}$ over the drift duration $\tau_{i}$, denoted by $\boldsymbol{x}_{ki}=\mathbb{E}[\boldsymbol{x}(t_{k};\tau_{i})]$ and $\Sigma_{ki}=\Sigma(t_{k};\tau_{i})=\mathbb{E}[\boldsymbol{x}(t_{k};\tau_{i})\boldsymbol{x}(t_{k};\tau_{i})^{\top}]$, can be written as follows:
@@
where $q(\cdot)$ is the inverse of the normal cumulative distribution function. Three uncertainty sources are considered in this case study: unmodeled acceleration, navigation uncertainty, and control execution error. $Q_{ki}$ is the discrete-time process noise covariance, representing the white-noise unmodeled acceleration. Eq. (6d) indicates that the covariance immediately after the loss of control has contributions from both navigation uncertainty and the control execution error. A simplified navigation model is introduced [18] to model the navigation covariance as a function of the range between the target and the servicer as follows:
@@ 6 Results and Analysis
Table 1: Domain Randomization used for the dataset generation. $\textsf{linspace}(x_{\text{min}},x_{\text{max}},n)$ represents an $n$-element vector of evenly spaced points between $x_{\text{min}}$ and $x_{\text{max}}$.
@@
$i$ 51.641[]51.641\ $^{\circ}$] $\textsf{linspace}(1\cdot 10^{-3},\pi/2,10)\ [\text{rad}]$
@@
$(r_{\text{KOZ}}\ [\text{km}],r_{\text{AE}}\ [\text{km}],t_{\text{switch}}[\text{min.}])$ ( $1.0,0.5,77$ ) ( $6.0,1.0,198$ )
@@
Figure 3: Sampled 100 trajectories from the dataset for two rendezvous scenarios. The lower panels provide zoomed-in views of the corresponding upper panels.
@@ 6.1 Scenarios and Benchmarks
The training dataset is constructed by solving a batch of single-objective SCP problems, each initialized from a convex solution with a selected time horizon $N$ that omits the passive safety constraint in Eq. (5d) and minimizes the total control cost. Table 1 summarizes the scenario setup and the domain randomization used to generate the dataset. Specifically, during dataset generation, the flight time is randomly drawn from a set of evenly spaced values between two extremes. In total, 80,000 problem instances are generated. For each instance, both the convex relaxation and the corresponding SCP solution initialized with the convex solution are computed and stored. Only those trajectories for which the SCP successfully converges are retained, ensuring that the dataset consists of an even distribution of convex–nonconvex solutions, where the nonconvex solution is feasible with respect to the original optimization problem. Additionally, during inference, the model initializes the RTG value using the objective of the convex relaxation, which is the lower bound under the time discretization based on the fixed time step $\Delta t$.
@@ 6.1.2 Far-range elliptic orbit rendezvous
(b) Generalization across dynamics and flight time with the elliptic orbit scenario. The test data is not seen by the model during the training, yet is sampled from the same distribution.
(c) Generalized across the dynamics and flight time with the elliptic orbit scenario. The test data is taken from SpaceTrack’s medium-sized debris catalog (cf. Figure 5).
Figure 4: Performance analysis of the SCP with ART initialization compared to that with CVX initialization, tested for 2,000 scenarios. The statistics are binned across the cumulative constraint violation of the CVX solution.
The performance of the ART is compared with the solution of the SCP using the convex solution for its initialization. For simplicity, the following notation is used in the remainder of the paper:
@@ 6.2 RQ1: Generalization across Flight Time and Dynamics
The generalizability of the ART inference is evaluated based on the quality of the generated trajectories as warm starts for the subsequent SCP. To this end, 2,000 withheld test scenarios are generated, and the performance of CVX-SCP and ART-SCP is compared across three statistical metrics: (i) the success rate of convergence in SCP, (ii) the optimality of the resulting objective values, and (iii) the number of iterations required for SCP convergence. It is important to note that the baseline CVX solutions display a wide variation in cumulative trajectory-level constraint violations (i.e., $c_{1}$), with some trajectories even satisfying the nonconvex constraints incidentally. These cumulative constraint violations provide an indicator of the difficulty in converging to feasible solutions. Accordingly, the problems are first grouped into bins based on the cumulative constraint violation of the CVX solutions, after which the statistical analyses are performed. Three experiments are conducted to evaluate ART’s generalizability, as summarized in Figure 4. For each experiment, results are presented in three forms: (i) histograms indicating whether SCP converged or not, (ii) violin plots of optimality, measured as the deviation from the CVX solution, and (iii) violin plots of the iteration counts required for SCP to converge. These visualizations are organized according to the binned groups of cumulative constraint violation in the CVX solutions.
@@
First, generalization over the flight time [17] is evaluated in the context of the LEO rendezvous scenario [18], formulated as a fixed-time, single-objective nonconvex OCP. Figure 4(a) presents the statistical analysis of the ART-generated trajectories. For this analysis, the test data consist of unseen scenarios sampled from the same distribution as the training dataset. The results show that ART-generated trajectories consistently outperform the baseline across all three metrics, demonstrating higher success rates, improved optimality, and reduced iteration counts to convergence. In particular, when cumulative constraint violations are low, the objective values at convergence are nearly identical; however, as the level of violation increases, the advantage of using ART-generated initial guesses becomes increasingly pronounced. This behavior occurs because initial guesses with low cumulative constraint violation require only minor adjustments to be rendered feasible, whereas those with high violations benefit from the substantially different trajectories proposed by ART, which guide the optimizer toward convergence within a more globally optimal basin in solution space. Overall, the results show that ART learns an effective control policy across varying flight times with reliability comparable to traditional convex-based methods. Moreover, it provides effective initial guesses that improve success rates, enhance optimality, and accelerate convergence in the subsequent SCP.
Next, generalization across both orbital domains and flight times is evaluated by varying the targets’ orbital elements, thereby inducing different underlying relative dynamics. The performance of the trained model under these conditions is summarized in Figure 4(b). Similar to the previous subsection, the test scenarios are drawn from the same distribution as the training dataset. While the success rate of convergence is slightly lower under ART initialization, both the achieved optimality and the number of iterations required for convergence remain equivalent or superior when using ART-generated initial guesses. This outcome reflects not only the inherent challenge of generalization but also the greater difficulty of the problem setting itself. Namely, far-range rendezvous (150–400 km separation), as opposed to the LEO case (<20<20 km), makes steering the system to the terminal state significantly more sensitive to small errors in the control inputs. Since ART inference entails sequential prediction of control inputs and forward dynamics propagation, the generated trajectories are dynamically feasible; however, strict satisfaction of the terminal constraint is not explicitly enforced, aside from conditioning on the queried goal token. In these more demanding scenarios, violations of the terminal constraint, which is an inherent limitation of the current ART inference architecture, exert a stronger influence on the subsequent SCP, often resulting in slower or incomplete convergence. Nevertheless, consistent with the earlier experiment on generalization across flight time, the relative advantage of ART becomes increasingly pronounced as the cumulative constraint violation of the CVX solution grows.
@@
As a demonstration of generalization, the model trained across varying flight times and underlying dynamics is further validated using the SpaceTrack orbital debris database [^3] https://www.space-track.org. Data acquisition date: Sep. 8, 2025. From a catalog of 5,000 medium-sized debris objects, 500 orbital elements are randomly selected within the ranges of semimajor axis (6,500<a<10,0006,500<a<10,000km), eccentricity ($1\cdot 10^{-3}<e<0.3$), and inclination ($0<i<90^{\circ}$). To promote diversity in the sampled orbital elements, targets are chosen in proportion to the logarithmic density across semimajor axis and eccentricity within each $(a,e)$ grid cell, thereby compensating for the strong concentration of debris in near-circular LEO. The resulting distribution of orbital elements for the sampled dataset is shown in Figure 5. Although the sampled orbital elements lie within the randomized training ranges, their values do not coincide with the discretized grid points, creating a more challenging test scenario for the model. The corresponding initial and boundary states are then generated by random sampling according to the specifications in Table 1.
@@
Figure 5: Distributions of the orbital elements of 500 sampled debris.
The results of the warm-starting performance evaluation are summarized in Figure 4(c). Favorable outcomes in terms of success rate, optimality, and iteration count are consistently confirmed across the cumulative constraint violation of the CVX solutions, particularly for cases with $c_{1}>5$. As noted earlier, the sampled scenarios are strongly biased toward near-circular, lower-altitude orbits, implying smaller initial separations and consequently lower sensitivity of the control. This observation supports the conclusion that the success rate of SCP warm-started with ART is significantly more advantageous in this test dataset, compared to the results shown in Figure 4(b).
@@ 6.3 RQ2: Prediction Capability of Mission Design Objectives
This subsection evaluates ART’s effectiveness as an efficient and accurate surrogate for the Pareto front, as well as its practical utility within the proposed framework. Representative outputs from the pipeline are shown in Figure 6. The input consists of three components: (i) target information (orbital elements), (ii) boundary conditions (initial and terminal), and (iii) the range of admissible flight times over which the tradespace is explored (cf. Figure 1). Based on this input, the pipeline autoregressively generates a batch (here, $B=6$) of near-optimal trajectories spanning different flight times (cf. Figure 6(a)) together with their corresponding tradespace in the multi-objective domain (cf. Figure 6(b)). In Figure 6(b), “SCP-min” denotes the minimum cost obtained across the converged solutions of CVX-SCP and ART-SCP, serving as an approximation of the closest achievable true Pareto set. These results demonstrate that ART consistently predicts the SCP cost more accurately than CVX. This predictive accuracy yields an offset that allows for a conservative approximation of the Pareto surface, inherently accounting for future constraint satisfaction—a feature that relaxed optimization formulations cannot capture by design.
@@
(c) Distribution of constraint violation and the runtime of each trajectory generation method. Each background color indicates fast and feasible (green), slow and feasible (yellow), fast and infeasible (blue), and slow and infeasible (red).
Figure 6: Near-Pareto-optimal trajectory generation via ART for a representative scenario.

### v2

@@ 5 Spacecraft Rendezvous in perturbed Keplerian orbits
This paper focuses on mission design for on-orbit servicing operations targeting multiple resident space objects in Earth orbit. In this scenario, a servicing spacecraft (servicer) executes an approach maneuver to a predefined waypoint in the vicinity of a target spacecraft (target). During the preliminary trajectory design phase of such missions, it is essential to conduct rapid trade studies between flight time and total control cost across a diverse set of candidate targets, each characterized by distinct orbital elements and, consequently, unique relative dynamics. In addition, trajectory design must account for the presence of uncertainty in the system. Accordingly, this paper considers optimization of a nominal deterministic trajectory that explicitly models uncertainty effects and satisfies mission constraints in a probabilistic sense.
@@ 5.1 Optimal Control Problem Formulation
The trajectory optimization problem described above can be formulated as a discrete-time MO-OCP for chance-constrained, fault-tolerant rendezvous trajectory design under uncertainty [57] as
@@
The state $\boldsymbol{x}\in\mathbb{R}^{6}$ is expressed in Relative Orbital Elements (ROE) [13][16]. Let the Keplerian orbital elements be $\boldsymbol{\alpha}_{c}=[a,e,i,\Omega,\omega,M]$, where $a$ is the semi-major axis, $e$ the eccentricity, $i$ the inclination, $\Omega$ the right ascension of the ascending node, $\omega$ the argument of periapsis, and $M$ the mean anomaly. The ROE are a particular set of differences between the orbital elements of the target and the servicer (also called the deputy), denoted as $\boldsymbol{\alpha}$ and $\boldsymbol{\alpha}_{d}$, respectively. The objective is to simultaneously minimize the total control cost, modeled as a sum of impulsive maneuver magnitudes, and the terminal time represented by the discrete horizon length $N$, leading to a flgiht time of $N\Delta t$. Eq. (5c) describes the discrete-time ROE dynamics with impulsive velocity change $\boldsymbol{u}\in\mathbb{R}^{3}$ applied in the Radial/Tangential/Normal (RTN) frame. The State Transition Matrix (STM) $\Phi(t_{k+1},t_{k})\in\mathbb{R}^{6\times 6}$ and control input matrix $\Gamma_{k}=\Gamma(t_{k})\in\mathbb{R}^{6\times 3}$ are expressed as an analytical function of the target’s orbital elements but not the relative states themselves [13][16], which offers a convenient expression as a linear time-varying system. Additionally, Eqs. (5e) and (5f) represent boundary conditions and maximum instantaneous control magnitude constraints, respectively. Finally, Eq. (5d) indicates the passive safety constraint, which ensures the deputy is outside of the ellipsoidal safety domains around the target not only for the entire duration of the controlled trajectory $t\in[0,t_{f}]$ but also for a predefined duration of the drift trajectories $\tau\in[0,\tau^{s}]$ after the sudden loss of control authority. The geometry of the safety domains at epoch $t_{k}$ is defined using the diagonal matrix $P_{k}\in\mathbb{R}^{6\times 6}$, which encodes the semi-axes of the ellipsoid within the RTN frame. In this case study, two safety domains are considered: the Approach Ellipsoid (AE) and the Keep-Out Zone (KOZ), both modeled as spheres. At the switching epoch, $t_{\text{switch}}$, the active safety domain transitions from the larger AE to the smaller KOZ, reflecting the practical scenario in which the allowable safety region contracts as the chaser spacecraft closes in on the target. Matrix $\Psi_{k}=\Psi(t_{k})\in\mathbb{R}^{6\times 6}$ denotes the first-order linear mapping between the ROE and the matrix expression of the relative position and velocity vector in the Cartesian coordinates within the RTN frame (denoted as $\boldsymbol{x}_{\text{RTN}}$) [13]. The passive safety constraint is formulated as a chance constraint to account for the inherent uncertainty that comes from navigation, control execution error, and unmodeled acceleration. At each epoch, the risk factor (i.e., failure probability) $\Delta_{k}$ is chosen as a parameter, and the probabilistic satisfaction of the constraint is enforced. Assuming Gaussian state uncertainty, the drifted mean state and covariance propagated from the time immediately after the control epoch $t_{k}$ over the drift duration $\tau_{i}$, denoted by $\boldsymbol{x}_{ki}=\mathbb{E}[\boldsymbol{x}(t_{k};\tau_{i})]$ and $\Sigma_{ki}=\Sigma(t_{k};\tau_{i})=\mathbb{E}[\boldsymbol{x}(t_{k};\tau_{i})\boldsymbol{x}(t_{k};\tau_{i})^{\top}]$, can be written as
@@
where $q(\cdot)$ is the inverse of the normal cumulative distribution function. Three uncertainty sources are considered in this case study: unmodeled acceleration, navigation uncertainty, and control execution error. $Q_{ki}$ is the discrete-time process noise covariance, representing the white-noise unmodeled acceleration. Eq. (6d) indicates that the covariance immediately after the loss of control has contributions from both navigation uncertainty and the control execution error. A simplified navigation model is introduced to model the navigation covariance as a function of the range between the target and the servicer as [57]
@@ 6 Results and Analysis
Table 1: Domain Randomization used for the training dataset generation. $\textsf{linspace}(x_{\text{min}},x_{\text{max}},n)$ represents an $n$-element vector of evenly spaced points between $x_{\text{min}}$ and $x_{\text{max}}$, including the two boundary values.
@@
$i$ 51.641[]51.641\ $^{\circ}$] $\textsf{linspace}(1\cdot 10^{-3},\pi-1\cdot 10^{-3},10)\ [\text{rad}]$
@@
$(r_{\text{AE}}\ [\text{km}],r_{\text{KOZ}}\ [\text{km}],t_{\text{switch}}[\text{min.}])$ ( $1.0,0.5,77$ ) ( $6.0,1.0,198$ )
@@
Figure 3: Sampled 100 trajectories from the dataset for two rendezvous scenarios. The lower panels provide zoomed-in views of the corresponding upper panels. Orange and red dotted circles indicate the two-dimensional projection of the AE and KOZ, respectively.
@@ 6.1 Scenarios and Benchmarks
The training dataset is constructed by solving a batch of single-objective SCP problems, each initialized from a convex solution with a selected time horizon $N$ that omits the passive safety constraint in Eq. (5d) and minimizes the total control cost. Table 1 summarizes the scenario setup and the domain randomization used to generate the training dataset. Specifically, during the dataset generation, the flight time is randomly drawn from a set of evenly spaced values between two extremes. In total, 80,000 problem instances are generated. For each instance, both the convex relaxation and the corresponding SCP solution initialized with the convex solution are computed and stored. Only those trajectories for which the SCP successfully converges are retained, ensuring that the dataset consists of an even distribution of convex–nonconvex solutions, where the nonconvex solution is feasible with respect to the original optimization problem. Additionally, during inference, the model initializes the RTG value using the objective of the convex relaxation, which is the lower bound under the time discretization based on the fixed time step $\Delta t$.
@@ 6.1.2 Far-range elliptic orbit rendezvous
(b) Generalization across dynamics and flight time with the elliptic orbit scenario. The test scenarios are not seen by the model during the training, yet are sampled from the same distribution.
(c) Generalization across the dynamics and flight time with the elliptic orbit scenario. The test scenarios are taken from SpaceTrack’s medium-sized debris catalog (cf. Figure 5).
Figure 4: Performance analysis of the SCP with ART initialization compared to that with CVX initialization, tested for 2,000 test scenarios. The statistics are binned across the cumulative constraint violation of the CVX solution.
The performance of ART is assessed by comparing the resulting SCP solutions initialized with ART-generated trajectories against those obtained using a conventional convex initialization. In particular, a distinct set of problem instances, separate from those used in the training dataset and referred to as test scenarios, is employed for the analysis. The design of the test scenarios for each analysis is described in the following sections. For simplicity, the following notation is used in the remainder of the paper:
@@ 6.2 RQ1: Generalization across Flight Time and Dynamics
The generalizability of the ART inference is evaluated based on the quality of the generated trajectories as warm starts for the subsequent SCP. To this end, 2,000 test scenarios are generated, and the performance of CVX-SCP and ART-SCP is compared across three statistical metrics: (i) the success rate of convergence in SCP, (ii) the optimality of the resulting objective values, and (iii) the number of iterations required for SCP convergence. It is important to note that the baseline CVX solutions display a wide variation in cumulative trajectory-level constraint violations (i.e., $c_{1}$), with some trajectories even satisfying the nonconvex constraints incidentally. These cumulative constraint violations provide an indicator of the difficulty in converging to feasible solutions. Accordingly, the problems are first grouped into bins based on the cumulative constraint violation of the CVX solutions, after which the statistical analyses are performed. Three experiments are conducted to evaluate ART’s generalizability, as summarized in Figure 4. For each experiment, results are presented in three forms: (i) histograms indicating whether the SCPs converged or not, (ii) violin plots of optimality, measured as the deviation from the CVX solution (lower bound), and (iii) violin plots of the iteration counts required for the SCP to converge. These visualizations are organized according to the binned groups of cumulative constraint violation in the CVX solutions.
@@
First, generalization over flight time [11] is evaluated in the context of the LEO rendezvous scenario [57]. Figure 4(a) presents the statistical analysis of the ART-generated trajectories. For this analysis, the test scenarios consist of previously unseen instances sampled from the same distribution as the training dataset. The results show that initialization with the ART-generated trajectories leads to a higher success rate of the convergence in the SCP across a wide range of constraint violations in the CVX solutions. Meanwhile, the optimality of the converged solution and the iteration counts to convergence are statistically improved or equivalent to the CVX-warmstarting. Overall, the results demonstrate that ART learns an effective control policy across varying flight times with reliability comparable to that of traditional convex-based methods. Moreover, it provides effective initial guesses that improve success rates, enhance optimality, and accelerate convergence in the subsequent SCP.
Next, generalization across both orbital domains and flight times is evaluated by varying the targets’ orbital elements, thereby inducing different underlying relative dynamics. The performance of the trained model under these conditions is summarized in Figure 4(b). The test scenarios are drawn from the same distribution as the training dataset. The results present a similar trend to that of the LEO rendezvous scenario, where the success rate of the SCP’s convergence using ART initialization outperforms the CVX initialization across most of the $c_{1}$ bins, while both the achieved optimality and the number of iterations required for convergence remain equivalent or superior when using ART-generated initial guesses.
@@
As a demonstration of generalization, the model trained with the elliptic orbit rendezvous dataset with varying flight times and underlying dynamics is further validated using the test scenarios based on the SpaceTrack orbital debris database [^2] https://www.space-track.org. Data acquisition date: Jan. 4, 2026. From a catalog of 5,000 medium-sized debris objects, 500 orbital elements are randomly selected within the ranges of semimajor axis ($6,500<a<10,000$ km) and eccentricity ($1\cdot 10^{-3}<e<0.3$). To promote diversity in the sampled orbital elements, targets are chosen in proportion to the logarithmic density across semimajor axis and eccentricity within each $(a,e)$ grid cell, thereby compensating for the strong concentration of debris in near-circular LEO. The resulting distribution of orbital elements of the target is shown in Figure 5. Although the sampled orbital elements lie within the randomized training ranges, their values do not coincide with the discretized grid points, creating more challenging test scenarios for the model. The corresponding initial and boundary states are then generated by random sampling according to the specifications in Table 1. Overall, a total of 2,000 test scenarios are generated and evaluated.
@@
Figure 5: Distributions of the orbital elements of 500 sampled debris used for the test scenarios.
The results of the warm-starting performance evaluation are summarized in Figure 4(c). Favorable results in terms of success rate, optimality, and iteration count are consistently observed across the cumulative constraint violation of the CVX solutions, particularly for cases with $c_{1}>5$. Although the training dataset may contain up to $10^{6}$ grid points of the chief orbital elements (cf. Table 1), the test scenarios considered in this experiment still require interpolation over a large six-dimensional continuous space. This observation supports the capability of the trained model to generalize to previously unseen parameter values within the range of the training dataset.
@@ 6.3 RQ2: Prediction Capability of Mission Design Objectives
This subsection evaluates ART’s effectiveness as an efficient and accurate surrogate for the Pareto front, as well as its practical utility within the proposed framework. Representative outputs from ART are shown in Figure 6. The input consists of three components: (i) target orbital elements, (ii) boundary conditions (initial and terminal), and (iii) the range of admissible flight times over which the tradespace is explored (cf. Figure 1). In particular, the figure summarizes the outputs of the following test scenario inputs, which are provided to a model trained for the far-range elliptic-orbit rendezvous case:
$\boldsymbol{\alpha}(t_{0})$ $=[a\ [\text{km}],e,i\ [\text{rad}],\Omega\ [\text{rad}],\omega\ [\text{rad}],M\ [\text{rad}]]$
$=[9.50\cdot 10^{3},0.195,0.629,0.0,4.40,2.51],$ (11a)
$a\boldsymbol{x}_{\text{e}}(t_{0})$ $=a[\delta a,\delta\lambda,\delta e_{x},\delta e_{y},\delta i_{x},\delta i_{y}]\ \text{[km]}$
$=[-15.1,-300,53.1,-19.6,3.05,-0.18],$ (11b)
$4.16$ $\leq t_{f}\leq 6.0\ \text{[h]}.$ (11c)
Based on this, the pipeline autoregressively generates a batch (here, $B=6$) of near-optimal trajectories spanning different flight times (cf. Figure 6(a)) together with their corresponding tradespace in the multi-objective domain (cf. Figure 6(b)). In Figure 6(b), “SCP-min” denotes the minimum cost obtained across the converged solutions of CVX-SCP and ART-SCP, serving as an approximation of the closest achievable true Pareto set. These results demonstrate that ART consistently predicts the SCP cost more accurately than CVX. This predictive accuracy yields an offset that allows for a conservative approximation of the Pareto surface, inherently accounting for future constraint satisfaction—a feature that relaxed optimization formulations cannot capture by design.
@@
(c) Distribution of constraint violation and the runtime of each trajectory generation method. Each background color qualitatively indicates fast and feasible (green), slow and feasible (yellow), fast and infeasible (blue), and slow and infeasible (red).
Figure 6: Near-Pareto-optimal trajectory generation via ART for a representative test scenario (cf. Eq. (11)).